HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH-RESOLUTION SPECTROSCOPY AND ANALYSIS OF THE 3 μm REGION OF C_{2}H_{4} IN NATURAL ISOTOPIC ABUNDANCE.

\underline{O. BEN FATHALLAH} , Université de Reims Champagne Ardenne, CNRS, GSMA UMR 7331, 51097 Reims, France, Laboratory of Spectroscopy and Molecular Dynamics, National Superior School of Engineering of Tunis, 5 Av Taha Hussein, 1008 Tunis, Tunisia; \underline{M. ROTGER} , \underline{M. REY}, Université de Reims Champagne Ardenne, CNRS, GSMA UMR 7331, 51097 Reims, France; \underline{H. AROUI}, Laboratory of Spectroscopy and Molecular Dynamics, National Superior School of Engineering of Tunis, 5 Av Taha Hussein, 1008 Tunis, Tunisia; \underline{V. BOUDON}, Lab. ICB, UMR 6303 CNRS/Univ. Bourgogne, 9, av. Alain Savary, B.P. 47870, F-21078 Dijon, France; \underline{J. VANDER AUWERA} , SQUARES, C.P. 160/09, Université Libre de

Bruxelles, 50 avenue F. D. Roosevelt, B-1050 Brussels, Belgium; \underline{M.A. LOROÑO GONZALEZ} , Oficina 410, Departamento Académico de

Fisicoquímica, Facultad de Química e Ingenieria Química, Universidad Nacional Mayor de San Marcos, Lima, Peru

Ethylene is a natural gaseous compound present in the atmosphere of the Earth but also observed in the atmospheres of outer solar system bodies such as Neptune and Titan^{1} ^{2}. Optical remote sensing of this species in these environments requires knowledge of spectroscopic parameters characterizing its spectrum. Such an information is compiled in databases,^{3}^{,}^{4} their improvement being a long-lasting, constant effort.

We present here a detailed analysis and modeling of the strongly absorbing ν_{9} and ν_{11} fundamental bands of C_{2}H_{4} observed in the 3 μm region. Due to the complexity of the observed spectrum, we have built a reasonably complex polyad scheme that includes some fundamental and combination bands previously analyzed using the tensorial formalism developed in Dijon for asymmetric-top molecules^{5}. A fourpolyad system has been used to analyze a region where mainly five rovibrational modes: ν_{9}, ν_{11}, ν_{2} + ν_{12}, ν_{2} + ν_{10}, and 2ν_{10} + ν_{12} of the ethylene emerge. A first frequency analysis has been performed, providing 3328 assignments and 87 adjusted parameters with a root mean square of 5.9 × 10^{−3} cm^{−1} and a standard deviation of 2.0 × 10^{−3} cm^{−1}. The line intensities were then fitted with a root mean square deviation of 1.9%, using parameters determined from the analysis of the line positions. The band intensities were also determined and compared with previous work. A new line list of calculated lines in the 2900-3300 cm^{−1} region will be proposed to

1G.S. Orton et al, Icarus, 70, 1-12, (1987)

2A. Coustenis et al, Icarus, 124, 54-76, (1989)

3I.E. Gordon et al, J. Quant. Spectrosc. Radiat. Transf., 277, 107949, (2022)

4T. Delahaye et al., J. Mol. Spectrosc, 380, 111510, (2021)

5J.P. Champion et al, Academic Press Inc., 339-422, (1992)

p-number: p031  Submitted on Sun, 04 Jun 2023 09:32:30 +0200

HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

be added to the ECasDa, HITRAN^{6}, and GEISA databases^{7}.

6I.E. Gordon et al, J. Quant. Spectrosc. Radiat. Transf., 277, 107949, (2022) 7T. Delahaye et al, J. Mol. Spectrosc, 380, 111510, (2021)

p-number: p031  Submitted on Sun, 04 Jun 2023 09:32:30 +0200